Explain What formula you have used to determine the charge to mass ratio of the electron and what is

the origin of it?


1
Expert's answer
2021-02-10T10:05:31-0500

Suppose the electron moves in electric "\mathbf{E}" and magnetic "\mathbf{B}" fields. The charge to mass ratio of the electron can be found from two formulas: second Newton's law:


"\mathbf{F} = m\mathbf{a}"

and Lorentz force law:


"\mathbf{F} = e(\mathbf{E} + \mathbf{v}\times \mathbf{B})"

where "\mathbf{v}" is the velocity of electron. Since forces are equal, obtain:


"\left( \dfrac{e}{m} \right)(\mathbf{E} + \mathbf{v}\times \mathbf{B}) = \mathbf{a}"

This is the most general expression to determine the charge to mass ratio of the electron.